1. Consider a voltaic cell in which the following reaction takes place in basic medium at 25℃.


2 NO3- (aq) + 3 S2- (aq) + 4 H2O 🡪 3 S (s) + 2 NO (g) + 8 OH- (aq)

  1. Calculate E◦
  2. Write the Nernst equation for the cell voltage E.
  3. Calculate E under the following conditions: PNO = 0.994 atm, pH =13.7, [S2-] = 0.154 M, [NO3-] = 0.472 M
1
Expert's answer
2021-02-11T04:29:08-0500

1. E⁰ = E"_1^\u2070" + E"_2^\u2070"

E⁰ = 0.78 + 0.407

E⁰ = 1.187V

2. E = E⁰ - "\frac{0.059}{n}"log"\frac{[OH^-]⁸(p_{NO})^2}{[NO_3^{-2}]^2[S^{2-}]^³}"


E = E⁰ - "\frac{0.059}{6}"log"\frac{[OH^-]⁸(p_{NO})^2}{[NO_3^{-2}]^2[S^{2-}]^³}"


3. pH = 13.7

pOH = 0.3

[OH-] = 10-0.3 = 0.5M

[NO"_3^{-2}" ] = 0.472M

[S2-] =0.154M

pNO = 0.994 atm

E = 1.187 - "\frac{0.059}{6}"log"\frac{(0.5)⁸(0.994)²}{(0.472)²(0.154)³}"


E = 1.187 + "\frac{0.059}{6}" ×11 log 8.18


E = 1.187 + "\frac{0.059}{6}" × 0.912


E = 1.187 + 0.098

E = 1.285 V